the probability of observing defective item in a lot is 0.01.suppose 25 items are inspected from the lot,what is the probability that not more than 3 defective are found, exactly 4 defective are found.(max 1024)


1
Expert's answer
2022-04-26T13:26:22-0400

Let "X=" the number of defective items: "X\sim Bin (n, p)."

Given "p=0.01, q=1-p=0.99, n=25."


"P(X\le3)=P(X=0)+P(X=1)"

"+P(X=2)+P(X=3)"

"=\dbinom{25}{0}(0.01)^0(0.99)^{25-0}+\dbinom{25}{1}(0.01)^1(0.99)^{25-1}"

"+\dbinom{25}{2}(0.01)^2(0.99)^{25-2}+\dbinom{25}{3}(0.01)^3(0.99)^{25-3}"

"=0.99989307347"

"P(X=4)=\dbinom{25}{4}(0.01)^4(0.99)^{25-4}"

"=0.00010243058"